OpenStudy (anonymous):

Solve for x: -2(x + 3) = -2(x + 1) - 4. 2 3 All real numbers No solution

jigglypuff314 (jigglypuff314):

-2(x + 3) = -2(x + 1) - 4 distribute -2x - 6 = -2x - 2 - 4 add like terms -2x - 6 = -2x - 6 add 2x to both sides +2x +2x -6 = -6 -6 does equal -6 therefore "all real numbers" but keep in mind that if it was something like 2 = 5 2 does not equal 5 so "no solution"

OpenStudy (anonymous):

wow thanks and I didn't even tag you madal

OpenStudy (anonymous):

oops I mean medal.

jigglypuff314 (jigglypuff314):

lol yeah ;) and you're welcome :P

OpenStudy (anonymous):

Ok this one I have to wright the answer and then I could answer the rest because this is hard. :P

OpenStudy (anonymous):

Solve for x: 3 over 4 x + 5 over 8 = 4x x = __________________ Write your answer as a fraction in simplest form. Use the "/" symbol for the fraction bar.

jigglypuff314 (jigglypuff314):

(3/4)x + (5/8) = 4x make your life easier, multiply everything by 8 6x + 5 = 32x then subtract both sides by 6x -6x -6x 5 = 26x divide both sides by 26

OpenStudy (anonymous):

it equals to 0.1923076923

jigglypuff314 (jigglypuff314):

mmm I think they want you to keep it as a fraction :)

OpenStudy (anonymous):

well then is it 5/26

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

wow aswesome thank you alot for your help I could take the other questions from here I could do the rest k.

jigglypuff314 (jigglypuff314):

you're welcome :) and good luck!

OpenStudy (anonymous):

K